OpenStudy (sharpsage):

What are the steps you need to do to verify a trigonometric identity? I'm not sure how to do it. 1 - 2cos^2x = 2sin^2x - 1

OpenStudy (sharpsage):

\[1-2\cos^2x=2\sin^2x-1\]

OpenStudy (australopithecus):

You just want to show that one side is equal to the other side, so pick a side then apply trig identities until it looks the same as the side you did not touch

OpenStudy (jdoe0001):

hmmm if you were to solve say \(\bf sin^2(\theta)+cos^2(\theta)=1\) for \(cos^2(\theta)\) what would you end up with?

OpenStudy (triciaal):

one approach use identities you know to express terms differently on one side then simplify until the result is the same as the other side using sin^2 + cos ^2 = 1 working with the left hand side LHS

OpenStudy (triciaal):